OpenStudy (anonymous):

4x+7y=-80 3x+5y=-58 elimination using multiplacation.

OpenStudy (anonymous):

First, lets cancel the x's, by doing this, we need to multiply the first equation by 3 and the second equation by -4. 4x+7y=-80 -----> Multiplied by 3 3x+5y=-58 -----> Multiplied by -4 Can you tell me what each equation will be now? :o)

OpenStudy (anonymous):

12x+28y=-80 -12x +20y=-58

OpenStudy (anonymous):

4x+7y=-80 -----> 12x+21y= -240 3x+5y=-58 ----> -12x-20y = 232 Now we're going to add like this: 12x+21y= -240 + -12x-20y = 232 --------------- y= -8 Notice how we multiplied the equations so that the value of x would cancel out so we could find the value of y. Do you see how we did that?

OpenStudy (anonymous):

yeah, and x, we do the opposite with. replacing y's with -8's.

OpenStudy (anonymous):

Right, so replace the y with -8 in the first original equation: 4x+7y=-80 So it'd be: 4x+7(-8)=-80 Now, solve for x. Do you know how? :o)

OpenStudy (anonymous):

x=6.

OpenStudy (anonymous):

Remeber, ALWAYS replace in the ORIGINAL equation or your answer wont be correct. :o) 4x-56=-80 4x=-24 x= -6 I think you forgot the negative sign. :o) But great work :o)

OpenStudy (anonymous):

THANK YOU. so much better than my teacher.

OpenStudy (anonymous):

Anytime :o) If you ever need any other help just holler. :o)